# Smoothed Circulant Embedding with Applications to Multilevel Monte Carlo Methods for PDEs with Random Coefficients

### v2

@@ 1 Introduction
In this paper, we focus on the circulant embedding technique for sampling from the random parameter. While this offers a discrete representation of a random field on a given mesh, it is both exact at these grid points, and computationally efficient. This is in comparison with the KL-expansion, which, once truncated, offers a continuous approximation whose rate of convergence is intrinsically linked to the decay of the eigenvalues of the covariance function’s kernel, which has a direct impact on the computational cost. The circulant embedding method has been integrated with MLMC and MLMQMC in e.g. [Graham et al., 2011, Kuo et al., 2017].
@@ 2.1 Circulant Embedding
where $C:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a suitably chosen function. For simplicity, we set the mean of $Z$ to 00 throughout this work, but we note that everything can be easily extended to non-zero mean $m(\mathbf{x}):D\rightarrow\mathbb{R}$.
@@ 2.2 Smooth Periodisation
There are instances when the circulant embedding matrix $S$ is not positive definite, depending generally on the covariance function used. As such, the Fourier diagonsalisation cannot be applied directly. To ensure the positive definiteness of $S$, Bachmayr et al. suggest in [Bachmayr et al., 2020] periodically extending the covariance function of the Gaussian field and padding the resulting covariance matrix before embedding it to obtain a circulant structure. Here, the analysis is developed for the case when the function $C$ is the Matérn covariance, given by:
where $\nu>0$ is the smoothness exponent, $\lambda>0$ denotes the length scale, $\sigma^{2}>0$ is the marginal variance and $K_{\nu}$ stands for the modified Bessel function of the second kind. Without loss of generality, we assume that the sampling domain $D$ is contained in $[0,1]^{d}$, so that $C:[-1,1]^{d}\rightarrow\mathbb{R}$.
@@ 2.3 Eigenvalue Decay
Throughout the remainder of the paper, let $\{\Lambda^{\text{ord}}_{j}\}_{j=1}^{s}$ be the circulant embedding eigenvalues $\{\Lambda_{j}\}_{j=1}^{s}$ arranged in non-increasing order.
A key factor in the study of the proposed smoothing technique is the decay rate of the ordered circulant embedding eigenvalues $\{\Lambda_{j}^{{\text{ord}}}\}_{j=1}^{s}$, obtained from the discrete Fourier transform as described in section 2.1. These are intricately linked with the eigenvalues of the covariance operator of the periodised covariance (2.6), which are given by its Fourier transform:
@@
In fact, the discrete eigenvalues $\{\Lambda_{j}\}_{j=1}^{s}$, when scaled by $m^{-d}$, can be seen as a rectangle rule approximation of the continuous eigenvalues $\lambda_{j}^{\text{ext}}$ above, with mesh size $m$, assuming the same discretisation parameter $m$ in all $d$ dimensions. For different mesh size $m_{i},i=1,\dotsc,d$, the scaling coefficient is $\prod_{i=1}^{d}m_{i}^{-1}$. Thus, as shown in Theorem 3.1 in Graham et al. [2018], the scaled discrete eigenvalues of the extended circulant embedding matrix of $C^{\text{ext}}$ converge to the continuous eigenvalues of the covariance operator of the periodised covariance $C^{\text{ext}}$.
In turn, the eigenvalues of the covariance operator of the periodised covariance are related to the exact Karhunen–Loève (KL) eigenvalues $\lambda_{j}$ of the operator of the original covariance with kernel $C$ defined on $L^{2}(D)$. The KL eigenvalues $\lambda_{j}$ emerge in the KL expansion of the continuous random field $Z$, given by:
where $\{\xi_{j}\}_{j\in\mathbb{N}}\sim\mathcal{N}(0,1)$ i.i.d., and $\{b_{j}\}_{j\in\mathbb{N}}$ are the $L^{2}(D)$-orthonormal eigenfunctions of the covariance operator. More specifically, Bachmayr et al. show in [Bachmayr et al., 2018, Section 3] that, when extending the domain $D\subset\mathbb{R}^{d}$ of a random field to a periodic domain $\mathbb{T}^{d}$, the eigenfunctions of the corresponding periodized covariance are explicitly known trigonometric functions (see [Bachmayr et al., 2018, Eq. (57)]), and the eigenvalues are Fourier coefficients of an even and $\mathbb{T}$-periodic function (see [Bachmayr et al., 2018, Eq. (59)]). By restricting back to the original domain $D$, we obtain a KL-type expansion of the random field on the original domain $D$. Exploiting this, Bachmayr et al. explain further in [Bachmayr et al., 2018] that the ordered eigenvalues $\{\lambda_{j}^{{\text{ext}}}\}_{j=1}^{s}$ of the periodised covariance $C^{\text{ext}}$ in the smooth periodisation case decay at precisely the same rate as the eigenvalues of the covariance operator of the original covariance $C$. This is proven for a general class of covariance functions in [Bachmayr et al., 2018, Eq. (64)], however here we focus on the Matérn case, stated in Lemma 1 below.
Finally, to find the decay rate of the discrete eigenvalues, we note that:
Then, as shown in the proof of Theorem 6 in [Bachmayr et al., 2020], the discrete eigenvalues converge to the continuous eigenvalues faster than $|\lambda^{\text{ext}}_{j}|$ decays. Hence, $\left\{\left(\prod_{i=1}^{d}m_{i}^{-1}\right)\Lambda^{\text{ord}}_{j}\right\}_{j=1}^{s}$ decay at the same rate as $\lambda_{j}^{\text{ext}}$, and, therefore, as $\lambda_{j}$.
In our numerical experiments, we also consider the separable exponential covariance function, given by:
with $p=1$. Here, $\|\cdot\|_{p}$ denotes the $\ell_{p}$ norm on $\mathbb{R}^{d}$, and the parameters $\sigma^{2}$ and $\lambda$ stand for the marginal variance and the correlation length of $C$, respectively. In this case, the eigenvalues in the circulant embedding factorisation are, in fact, real and positive, so that no padding is necessary (this is shown in Lemma 2 below).
Thus, the following two lemmas give the decay rate of the ordered eigenvalues $\{\Lambda_{j}^{\text{ord}}\}_{j=1}^{s}$ for the case of Matérn and separable exponential covariance functions, respectively.
@@
Let $C$ be the Matèrn covariance (2.5), and suppose $\nu,\lambda$ and $\kappa$ satisfy Eq. (2.7). Let $\{\Lambda_{j}^{{\text{ord}}}\}_{j=1}^{s}$ denote the ordered eigenvalues of the (extended) embedding matrix $S\in\mathbb{R}^{s\times s}$, with $s=\prod_{i=1}^{d}2(m_{i}+J_{i})$. Then for $j=1,\dots,s$:
@@
This follows from the fact that the discrete eigenvalues decay at the same rate as the continuous eigenvalues, as shown above. The decay rate for the continuous eigenvalues is given in [Graham et al., 2015, Corollary 5]. ∎
@@
$\left(\prod_{i=1}^{d}m_{i}^{-1}\right)\,\Lambda_{j}^{{\mathrm{ord}}}\lesssim\begin{cases}j^{-2},&d=1\\ j^{-2+\varepsilon},\varepsilon>0,&d>1.\end{cases}$ (2.12)
@@
First, Eq. (2.11) is a direct application of Theorem 2 in [Dietrich and Newsam, 1997]. Further, this implies that no padding is necessary, as the embedding matrix is symmetric positive definite.
@@
$\mathbf{s}^{nD}=\frac{1}{\sigma^{n}}\mathbf{s}^{1D}\otimes\dotsc\otimes\mathbf{s}^{1D},$
where $\mathbf{s}^{nD}$ is the first column of the $n$-dimensional embedding matrix $S$, and $\otimes$ denotes the Kronecker product. Further, the discrete Fourier matrix satisfies [Cooley et al., 1969]:
$F^{nD}=F^{1D}\otimes\dotsc\otimes F^{1D},$
@@
$\boldsymbol{\Lambda}^{nD}=\boldsymbol{\Lambda}^{1D}\otimes\dotsc\otimes\boldsymbol{\Lambda}^{1D}.$ (2.13)
To obtain the decay rate of the eigenvalues in the $n$-dimensional case, we derive the two-dimensional case following the argument in [Charrier, 2012, Section 7.1]; a similar argument holds in any dimension $n$. To this end, let $\{\Lambda^{1D}_{j}\}_{j\geq 1}$ and $\{\Lambda^{2D}_{j}\}_{j\geq 1}$ denote the eigenvalues sorted in decreasing order corresponding to the one- and two-dimensional case, respectively. Then, for any $j\geq 1$, there exists $(i,k)$ such that $\Lambda_{j}^{2D}=\Lambda_{i}^{1D}\Lambda_{k}^{1D}$ from Eq. (2.13). In fact, there might be more than one pair $(i,k)$ that gives the eigenvalue $\Lambda_{j}^{2D}$. To ensure uniqueness, we assume that any such pairs $(i,k)$ are further ordered lexicographically. Using this, we can then define the following bijective function:
$g:$ $(\mathbb{N}^{*})^{2}\rightarrow\mathbb{N}^{*},$
$(i,k)\mapsto j\text{ such that }\Lambda_{j}^{2D}=\Lambda_{i}^{1D}\Lambda_{k}^{1D}.$
Take $(i,k)$ such that $ik=j$. Then if $(p,q)$ is such that $\Lambda_{p}^{1D}\Lambda_{q}^{1D}\geq\Lambda_{i}^{1D}\Lambda_{k}^{1D}$, we have:
$\left(\frac{8\sigma^{2}}{\lambda\pi^{2}}\right)^{2}\frac{1}{p^{2}q^{2}}\geq\Lambda^{1D}_{p}\Lambda^{1D}_{q}\geq\Lambda^{1D}_{i}\Lambda_{k}^{1D}\geq\left(\frac{2\lambda\sigma^{2}}{1+4\lambda^{2}\pi^{2}}\right)^{2}\frac{1}{j^{2}},$
using the bound (see [Charrier, 2012, Section 7.1] for details):
$\frac{2\lambda\sigma^{2}}{1+4\lambda^{2}\pi^{2}}\frac{1}{j^{2}}\leq\Lambda^{1D}_{j}\leq\frac{8\sigma^{2}}{\lambda\pi^{2}}\frac{1}{j^{2}}.$
This yields the following bound for $g(i,k)$:
$g(i,k)$ $\leq Card\{(p,q)|\Lambda^{1D}_{p}\Lambda^{1D}_{q}\geq\Lambda_{i}^{1D}\Lambda^{1D}_{k}\}$
$\leq Card\left\{(p,q)|pq\leq\frac{4(1+4\pi^{2}\lambda^{2})}{\lambda^{2}\pi^{2}}j\right\}$
$\leq C^{\prime}_{\lambda}j\log(j),$
where $C^{\prime}_{\lambda}$ is a constant that only depends on the correlation length $\lambda$. Then we can conclude that for any $\varepsilon>0$, there exists a constant $C_{\lambda,\varepsilon}$ such that for all $i,k\geq 1$:
$g(i,k)\leq C_{\lambda,\varepsilon}(ik)^{1+\varepsilon}.$
Thus, the eigenvalues satisfy:
$\Lambda^{2D}_{j}=\Lambda^{1D}_{g(i,k)}=\Lambda^{1D}_{i}\Lambda^{1D}_{k}\leq C\frac{1}{(ik)^{2}}\leq C_{\lambda,\varepsilon}\frac{1}{g(i,k)^{2-\varepsilon}}=C_{\lambda,\varepsilon}\frac{1}{j^{2-\varepsilon}},$
for any $\varepsilon>0$.
Hence, the eigenvalues corresponding to the $n$-dimensional case decay at the rate:
$\left(\prod_{i=1}^{d}m_{i}^{-1}\right)\Lambda^{nD,{{\mathrm{ord}}}}_{j}\lesssim j^{-2+{\varepsilon}},{\text{ for any }\varepsilon>0.}$
@@ 2.4 Smoothing
The choice of norm for the smoothing error $\mathbb{E}\left[\|\mathbf{Z}-\tilde{\mathbf{Z}}\|\right]$ is application dependent. Throughout the paper, we use the $L^{\infty}$ norm, as this is the natural norm that emerges in the finite element discretisation error analysis for the application at hand, as shown in Theorem 3 and Corollary 1. For different applications, one can adapt these results using the equivalence of norms.
@@ 3.6 MC and MLMC with Smoothed Circulant Embedding
$\mathbb{E}\left[|Q_{h}-\tilde{Q}_{h}|^{p}\right]\lesssim s^{-\frac{p}{2}}\left(\Lambda_{{s-{\tau}+1}}^{{\mathrm{ord}}}\right)^{\frac{p}{2}}\,{\tau}^{p}.$
@@
$\mathbb{E}\left[|Q_{h}-\tilde{Q}_{h}|^{p}\right]\lesssim(s-{\tau}{+1})^{-p}\,{\tau}^{p}.$
@@
$\mathbb{E}\left[|Q_{h}-\tilde{Q}_{h}|^{p}\right]\lesssim s^{-\frac{p}{2}}{\left(\prod_{i=1}^{d}m_{i}\right)^{\frac{p}{2}}}(s-{\tau}{+1})^{-\frac{p}{2}(1+\frac{2\nu}{d})}\,{\tau}^{p}.$
@@
Note that here we are computing a permutation $\pi:\{1,2,\dotsc,s_{\ell}\}\rightarrow\{1,2,\dotsc,s_{\ell}\}$ such that the ordered eigenvalues $\{\Lambda_{j}^{\mathrm{ord}}\}_{j=1}^{s_{\ell}}$ are essentially $\{\Lambda_{\pi(j)}\}_{j=1}^{s_{\ell}}$. The vector $\tilde{\boldsymbol{\Lambda}}$ is then obtained by applying the inverse permutation $\pi^{-1}$ to the truncated eigenvalues.

### v3

@@ 1 Introduction
In this paper, we focus on the circulant embedding technique for sampling from the random parameter. While this offers a discrete representation of a random field on a given mesh, it is both exact at these grid points, and computationally efficient. This is in comparison with the KL-expansion, which, once truncated, offers a continuous approximation whose rate of convergence is intrinsically linked to the decay of the eigenvalues of the covariance function’s kernel, which has a direct impact on the computational cost. Nevertheless, the KL-expansion and the circulant embedding method share certain analytical properties (as discussed further in section 2.3). The circulant embedding method has been integrated with MLMC and MLMQMC in e.g. [Graham et al., 2011, Kuo et al., 2017].
@@ 2.1 Circulant Embedding
where $C:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a suitably chosen function. For simplicity, we set the mean of $Z$ to 00 throughout this work, but we note that everything can be easily extended to non-zero mean $m(\mathbf{x}):D\rightarrow\mathbb{R}$. Although the circulant embedding methodology is generally applicable, in the analysis and numerical examples presented in this work we focus on using either a Matérn or separable exponential covariance function. The Matérn covariance function is given by:
where $\nu>0$ is the smoothness exponent, $\lambda>0$ denotes the length scale, $\sigma^{2}>0$ is the marginal variance and $K_{\nu}$ stands for the modified Bessel function of the second kind. The separable exponential covariance function, on the other hand, is given by:
with $p=1$. Here, $\|\cdot\|_{p}$ denotes the $\ell_{p}$ norm on $\mathbb{R}^{d}$, and the parameters $\sigma^{2}>0$ and $\lambda>0$ stand for the marginal variance and the length scale of $C$, respectively.
@@ 2.2 Smooth Periodisation
There are instances when the circulant embedding matrix $S$ is not positive definite, depending generally on the covariance function used. As such, the Fourier diagonsalisation cannot be applied directly. To ensure the positive definiteness of $S$, Bachmayr et al. suggest in [Bachmayr et al., 2020] periodically extending the covariance function of the Gaussian field and padding the resulting covariance matrix before embedding it to obtain a circulant structure. Here, the analysis is developed for the case when the function $C$ is the Matérn covariance, given by (2.2). Without loss of generality, we assume that the sampling domain $D$ is contained in $[0,1]^{d}$, so that $C:[-1,1]^{d}\rightarrow\mathbb{R}$. Note that for the separable exponential covariance function, given by (2.3), the eigenvalues of $S$ are, in fact, real and positive, so that no padding is necessary.
@@ 2.3 Eigenvalue Decay
Throughout the remainder of the paper, let $\{\Lambda^{\text{ord}}_{j}\}_{j=1}^{s}$ be the circulant embedding eigenvalues $\{\Lambda_{j}\}_{j=1}^{s}$ arranged in non-increasing order. The main results in this section, given in Lemma 1 and 2 below, provide a decay rate (in $j$) of $\{\Lambda^{\text{ord}}_{j}\}_{j=1}^{s}$. This will be a key factor in the study of the proposed smoothing technique. The proofs of Lemma 1 and 2 are based on the following two main ideas.
Firstly, as shown in Theorem 3.1 in Graham et al. [2018], the scaled (discrete) eigenvalues of the extended circulant embedding matrix of the periodised covariance $C^{\text{ext}}$ converge to the (continuous) eigenvalues of the covariance operator of $C^{\text{ext}}$ as the mesh sizes $h_{i},i=1,\dots,d$, tend to zero. This can be seen by comparing the expression for the discrete eigenvalues $\{\Lambda_{j}\}_{j=1}^{s}$, obtained from the discrete Fourier transform as described in section 2.1, with that of the continuous eigenvalues, given by the Fourier transform of $C^{\text{ext}}$:
@@
The discrete eigenvalues, when scaled by $\prod_{i=1}^{d}m_{i}^{-1}$, can thus be seen as a rectangle rule approximation of the continuous eigenvalues.
Secondly, it is shown in [Bachmayr et al., 2018] that the ordered eigenvalues $\{\lambda_{j}^{{\text{ext}}}\}_{j=1}^{s}$ of the periodised covariance $C^{\text{ext}}$ in the smooth periodisation case decay at precisely the same rate as the eigenvalues $\{\lambda_{j}\}_{j=1}^{s}$ of the covariance operator of the original covariance $C$. This is proven for a general class of covariance functions in [Bachmayr et al., 2018, Eq. (64) and (65)], however here we focus on the Matérn case, stated in Lemma 1 below. More details on this are given in Remark 1.
Finally, the following two lemmas give the decay rate of the ordered eigenvalues $\{\Lambda_{j}^{\text{ord}}\}_{j=1}^{s}$ for the case of Matérn and separable exponential covariance functions, respectively. Following from the observations above, the decay rate we obtain for the discrete eigenvalues $\{\Lambda^{\text{ord}}_{j}\}_{j=1}^{s}$ is the same as that of the continuous eigenvalues $\{\lambda_{j}\}_{j=1}^{\infty}$ of the covariance operator of the original covariance $C$.
@@
Let $C$ be the Matèrn covariance (2.2). Suppose the padding parameters $J_{i},i=1,\dotsc,d$, are chosen such that $\kappa$ satisfies Eq. (2.8). Let $\{\Lambda_{j}^{{\text{ord}}}\}_{j=1}^{s}$ denote the ordered eigenvalues of the (extended) embedding matrix $S\in\mathbb{R}^{s\times s}$, with $s=2^{d}\prod_{i=1}^{d}(m_{i}+J_{i})$. Then for $j=1,\dots,s$:
@@
To find the decay rate of the discrete eigenvalues, we note that:
Then, because the smoothed periodisation preserves the regularity of the Matérn covariance function by construction, we can use Lemma 1 in Bachmayr et al. [2020] to infer that the discrete eigenvalues converge to the continuous eigenvalues faster than $|\lambda^{\text{ext}}_{j}|$ decays. Hence, $\left\{\left(\prod_{i=1}^{d}m_{i}^{-1}\right)\Lambda^{\text{ord}}_{j}\right\}_{j=1}^{s}$ decay at the same rate as $\lambda_{j}^{\text{ext}}$, and, therefore, as $\lambda_{j}$ by [Bachmayr et al., 2018, Eq. (64) and (65)]. The decay rate for the continuous eigenvalues is given in [Graham et al., 2015, Corollary 5], which completes the proof. ∎
@@
$\left(\prod_{i=1}^{d}m_{i}^{-1}\right)\,\Lambda_{j}^{{\mathrm{ord}}}\lesssim{j^{-2}(\log j)^{2(d-1)},\quad\forall d\geq 1.}$ (2.11)
@@
First, Eq. (2.10) is a direct application of Theorem 2 in [Dietrich and Newsam, 1997]. This implies that no padding is necessary, as the embedding matrix is symmetric positive definite.
@@
$\mathbf{s}^{dD}=\frac{1}{\sigma^{n}}\mathbf{s}^{1D}\otimes\dotsc\otimes\mathbf{s}^{1D},$
where $\mathbf{s}^{dD}$ is the first column of the $d$-dimensional embedding matrix $S$, and $\otimes$ denotes the Kronecker product. Further, the discrete Fourier matrix satisfies [Cooley et al., 1969]:
$F^{ndD}=F^{1D}\otimes\dotsc\otimes F^{1D},$
@@
$\boldsymbol{\Lambda}^{dD}=\boldsymbol{\Lambda}^{1D}\otimes\dotsc\otimes\boldsymbol{\Lambda}^{1D}.$ (2.12)
We then use Theorem 1 in Krieg [2018] together with the decay rate of the one-dimensional eigenvalues above to infer that:
$\left(\prod_{i=1}^{d}m_{i}^{-1}\right)\Lambda^{dD,{{\mathrm{ord}}}}_{j}\lesssim j^{-2}(\log j)^{2(d-1)},$
with a constant that depends on the dimension $d$.
This completes the proof. ∎
Remark 1.
As the results in this section show, there is an intricate link between the circulant embedding method and the KL expansion: the discrete and continuous eigenvalues decay at the same rate, and the (scaled) discrete eigenvalues converge to the continuous eigenvalues. Furthermore, consider the KL expansion of the continuous random field $Z$, given by:
where $\{\xi_{j}\}_{j\in\mathbb{N}}\sim\mathcal{N}(0,1)$ i.i.d., and $\{\lambda_{j}\}_{j\in\mathbb{N}}$ and $\{b_{j}\}_{j\in\mathbb{N}}$ are the eigenvalues and $L^{2}(D)$-orthonormal eigenfunctions of the covariance operator, respectively. Bachmayr et al. show in [Bachmayr et al., 2018, Section 3] that, when extending the domain $D\subset\mathbb{R}^{d}$ of a random field to a periodic domain $\mathbb{T}^{d}$, the eigenfunctions of the corresponding periodized covariance are explicitly known trigonometric functions (see [Bachmayr et al., 2018, Eq. (57)]), and the eigenvalues are Fourier coefficients of an even and $\mathbb{T}$-periodic function (see [Bachmayr et al., 2018, Eq. (59)]). By restricting back to the original domain $D$, we obtain an alternative series (KL-type) expansion of the random field on the original domain $D$. Evaluated on a grid, the expressions for the discrete random field $\mathbf{Z}$ obtained from the circulant embedding method and this KL-type expansion coincide due to aliasing effects. Note however, that the smoothing technique introduced in this work and the truncation of the KL-type expansion (similar to Teckentrup et al. [2013]) are not equivalent.
@@ 2.4 Smoothing
The choice of norm for the smoothing error $\mathbb{E}\left[\|\mathbf{Z}-\tilde{\mathbf{Z}}\|\right]$ is application dependent. Throughout the paper, we use the $L^{\infty}$ norm, as this is the natural norm that emerges in the finite element discretisation error analysis for the application at hand, as shown in Theorem 3 and Corollary 1. For different applications, one can adapt these results using the equivalence of norms. However, this will lead to different convergence rates due to the presence of s-dependent factors.
@@ 3.6 MC and MLMC with Smoothed Circulant Embedding
$\mathbb{E}\left[|Q_{h}-\tilde{Q}_{h}|^{p}\right]\lesssim s^{-\frac{p}{2}}\left(\Lambda_{{s-{\tau}+1}}^{{\mathrm{ord}}}\right)^{\frac{p}{2}}\,{\tau}^{p},$
where $s$ if the size of the embedding matrix after padding, so that $s=2^{d}\prod_{i=1}^{d}(m_{i}+J_{i})$, for discretisation parameters $m_{i},i=1,\dotsc,d$ and padding values $J_{i},i=1,\dotsc,d$.
@@
$\mathbb{E}\left[|Q_{h}-\tilde{Q}_{h}|^{p}\right]\lesssim(s-{\tau}{+1})^{-p}\,{\tau}^{p},$
where $s$ if the size of the embedding matrix, so that $s=2^{d}\prod_{i=1}^{d}m_{i}$, for discretisation parameters $m_{i},i=1,\dotsc,d$.
@@
$\mathbb{E}\left[|Q_{h}-\tilde{Q}_{h}|^{p}\right]\lesssim s^{-\frac{p}{2}}{\left(\prod_{i=1}^{d}m_{i}\right)^{\frac{p}{2}}}(s-{\tau}{+1})^{-\frac{p}{2}(1+\frac{2\nu}{d})}\,{\tau}^{p},$
where $s$ if the size of the embedding matrix after padding, so that $s=2^{d}\prod_{i=1}^{d}(m_{i}+J_{i})$, for discretisation parameters $m_{i},i=1,\dotsc,d$ and padding values $J_{i},i=1,\dotsc,d$.
@@
Note that here we are computing a permutation $\pi:\{1,2,\dotsc,s_{\ell}\}\rightarrow\{1,2,\dotsc,s_{\ell}\}$ such that the ordered eigenvalues $\{\Lambda_{j}^{\mathrm{ord}}\}_{j=1}^{s_{\ell}}$ are given by $\{\Lambda_{\pi(j)}\}_{j=1}^{s_{\ell}}$. The vector $\tilde{\boldsymbol{\Lambda}}$ is then obtained by applying the inverse permutation $\pi^{-1}$ to the truncated eigenvalues.